# Abelian congruences and similarity in varieties with a weak difference term

### v1

@@ 8. Examples
    the division ring associated to ${\mathbf{A}},\mu$ and a transversal for $\mu$ can be any desired finite field $\mathbb{F}$;
@@
    The vector spaces ${}_{\overline{\mathbb{F}}}\operatorname{Grp}_{D({\mathbf{A}},\mu)}(\varphi,0_{C})$ as $C$ ranges over the $\alpha$-classes can have any finite dimensions.
@@
    The ranges of the $\mu$-classes within a given $\alpha$-class $C$ can be essentially any collection of subspaces of ${}_{\overline{\mathbb{F}}}\operatorname{Grp}_{{\mathbf{D}}({\mathbf{A}},\mu)}(\varphi,0_{C})$ including $0_{C}/\varphi$ itself.
Our recipe is an adaptation of a known construction [22] (see also [3]). We first give a useful method for constructing varieties with a prescribed weak difference term. For this, we need the following result characterizing weak difference terms at the level of varieties. Following [19], given a variety $\mathcal{V}$, let ${\mathbf{F}}={\mathbf{F}}_{\mathcal{V}}(3)$ be the $\mathcal{V}$-free algebra freely generated by $\{x,y,z\}$, let $\alpha=\operatorname{Cg}(x,z)$, $\beta=\operatorname{Cg}(x,y)$ and $\gamma=\operatorname{Cg}(y,z)$ calculated in ${\mathbf{F}}$, and define $\beta_{0}=\beta$, $\gamma_{0}=\gamma$, $\beta_{n+1}=\beta\vee(\alpha\wedge\gamma_{n})$, and $\gamma_{n+1}=\gamma\vee(\alpha\wedge\beta_{n})$ for $n\geq 0$.
Theorem 8.1.
@@
(1) $\Rightarrow$ (3) can be extracted from the proof of Theorem 4.8 (1) $\Rightarrow$ (2) $\Rightarrow$ (3) in [19]. Standard arguments with roots in [29, 34] can produce sets $\Sigma_{0},\Sigma_{1},\Sigma_{2},\ldots$ of identities witnessing (2) $\Leftrightarrow$ (3). To prove (3) $\Rightarrow$ (1), assume that $(x,d)\in\gamma_{m}$ and $(d,z)\in\beta_{m}$. Then by [35, Theorem 2.1 (3) $\Rightarrow$ (4)], there exist idempotent ternary terms $f_{1},\ldots,f_{n},g_{1},\ldots,g_{n}$ (for some $n$) such that $\mathcal{V}\models f_{i}(x,y,x)\approx g_{i}(x,y,x)$ for each $i\in[n]$, and for all ${\mathbf{A}}\in\mathcal{V}$ and all $a,b\in A$, if
@@
then $d(a,a,b)=b=d(b,a,a)$ and $d(b,b,a)=a=d(a,b,b)$. Now one can easily adapt the proof of [14, Theorem 1.2 (3) $\Rightarrow$ (1)] to get that $d$ is a weak difference term for $\mathcal{V}$. ∎
Definition 8.2.
A semilattice-over-Maltsev operation is any operation $d:A^{3}\rightarrow A$ constructed in the following way. Let $(S,\wedge)$ be a (meet) semilattice. For each $s\in S$ let $V_{s}$ be a non-empty set, and assume that $V_{s}\cap V_{t}=\varnothing$ for all $s,t\in S$ with $s\neq t$. For each pair $s,t\in S$ with $s\geq t$, assign a function $f_{(s,t)}:V_{s}\rightarrow V_{t}$, so that $f_{(s,s)}=\mathrm{id}_{V_{s}}$ and $f_{(s,u)}=f_{(t,u)}\circ f_{(s,t)}$ for all $s,t,u\in S$ with $s\geq t\geq u$. Finally, for each $s\in S$ let $m_{s}:(V_{s})^{3}\rightarrow V_{s}$ be a Maltsev operation.
@@
The identities (8.2) and (8.3) therefore hold in $\mathsf{HSP}({\mathbf{A}})$. Let ${\mathbf{F}}={\mathbf{F}}_{\mathsf{HSP}({\mathbf{A}})}(3)$ be the free algebra in $\mathsf{HSP}({\mathbf{A}})$ freely generated by $\{x,y,z\}$. As in the discussion preceding Theorem 8.1, let $\alpha=\operatorname{Cg}(x,z)$, $\beta=\operatorname{Cg}(x,y)$ and $\gamma=\operatorname{Cg}(y,z)$ be calculated in ${\mathbf{F}}$. Interpreting $d$ and $f_{0},\ldots,f_{3},g_{0},\ldots,g_{3}$ as elements of ${\mathbf{F}}$ in the usual way, we find that the identities (8.2) and (8.3) insure that $f_{0},\ldots,f_{3},g_{0},\ldots,g_{3}$ witness
@@
Thus $d$ is a weak difference term for $\mathsf{HSP}({\mathbf{A}})$ by Theorem 8.1. ∎
@@
Choose vector spaces $V_{(\ell,i)}$ over $\mathbb{F}$ and isomorphisms $\sigma_{(\ell,i)}:V_{(\ell,i)}\cong W_{(\ell,i)}$ for $\ell\leq m$ and $1\leq i\leq n_{\ell}$, so that the spaces $V_{(\ell,i)}$ ($0\leq\ell\leq m$, $0\leq i\leq n_{\ell}$) are pairwise disjoint. Also set $\sigma_{(\ell,0)}=\mathrm{id}_{V_{(\ell,0)}}$ for each $\ell\leq m$. Let $C_{\ell}$ be the (disjoint) union of $V_{(\ell,0)},V_{(\ell,1)},\ldots,V_{(\ell,n_{\ell})}$, and let $A$ be the (disjoint) union of $C_{0},\ldots,C_{m}$. Then let $\mu$ and $\alpha$ be the equivalence relations on $A$ whose equivalence classes are the spaces $V_{(\ell,i)}$ ($\ell\leq m$, $i\leq n_{\ell}$) and the sets $C_{0},\ldots,C_{m}$ respectively. Also let $B=V_{(0,0)}\cup\cdots\cup V_{(m,0)}$ and let $\psi$ be the equivalence relation on $B$ whose equivalence classes are $V_{(0,0)},\ldots,V_{(m,0)}$. We will build an algebra ${\mathbf{A}}$ with universe $A$ such that:
@@
  • Letting $\Delta=\Delta_{\mu,\alpha}$ and $\varphi=\overline{\alpha}/\Delta$, there is a bijection from $D({\mathbf{A}})$ to $B$ which sends $\varphi$ to $\psi$ and sends each $\varphi$-class $(\mu\cap(C_{\ell})^{2})/\Delta$ to $V_{(\ell,0)}$.
  • Modulo this bijection, for each $\alpha$-class $C_{\ell}$, the ranges (see Definition 5.5) of the $\mu$-classes $V_{(\ell,i)}$ contained in $C_{\ell}$ correspond precisely to the subspaces $W_{(\ell,i)}$ in $V_{(\ell,0)}$.
  • The division ring associated to ${\mathbf{A}},\mu$, and any transversal for $\mu$ is isomorphic to $\mathbb{F}$.
@@
$(\ell,i)\wedge(k,j)=\left\{\begin{array}[]{cl}(\ell,i)&\mbox{if $(\ell,i)=(k,j)$}\\ (\ell,0)&\mbox{if $k=\ell$ but $i\neq j$}\\ (0,0)&\mbox{if $k\neq\ell$.}\end{array}\right.$
@@
Now let $d:A^{3}\rightarrow A$ be the semilattice-over-Maltsev operation (see Definition 8.2) defined by the data $(S,\wedge)$, $(V_{s}:s\in S)$, $(f_{(s,t)}:s,t\in S,\,s\geq t)$, and the family of Maltsev operations $m_{s}(x,y,z)=x-y+z$ evaluated in each vector space $V_{s}$. That is, for all $a\in V_{s_{1}}$, $b\in V_{s_{2}}$ and $c\in V_{s_{3}}$, letting $t=s_{1}\wedge s_{2}\wedge s_{3}$ we have
[…]
For each $\ell\leq m$ define $\sigma_{\ell}:C_{\ell}\rightarrow V_{(\ell,0)}$ so that $\sigma_{\ell}{\upharpoonright}_{V_{(\ell,i)}}=\sigma_{(\ell,i)}$ for all $i\leq n_{\ell}$. For each $s\in S$ let $0_{s}$ denote the zero element of $V_{s}$. We define some unary operations on $A$.
@@
    $F_{(\ell,i)}^{g}(a)=\left\{\begin{array}[]{cl}(g\circ\sigma_{\ell})(a)&\mbox{if $x\in C_{\ell}$}\\ 0_{(k,0)}&\mbox{if $a\in C_{k}$ with $k\neq\ell$.}\end{array}\right.$
@@
    $G_{\ell}(a)=\left\{\begin{array}[]{cl}(g_{\ell}\circ\sigma_{0})(a)&\mbox{if $a\in C_{0}$}\\ 0_{(k,0)}&\mbox{if $a\in C_{k}$ with $k\neq 0$}\end{array}\right.$
@@
    $G_{\ell}^{\prime}(a)=\left\{\begin{array}[]{cl}(h_{\ell}\circ\sigma_{\ell})(a)&\mbox{if $a\in C_{\ell}$}\\ 0_{(k,0)}&\mbox{if $a\in C_{k}$ with $k\neq\ell$}\end{array}\right.$

### v2

@@ 8. Conclusion and questions
    For all $(\lambda_{i})_{i\in I}\in P$, the following are equivalent:
    1. (a)
      $\Psi((\lambda_{i})_{i\in I})\in\operatorname{End}({}_{\mathcal{R}}V)$.
    2. (b)
      For all $i,j\in I$ and all $r\in R_{ij}$, $\lambda_{i}\circ r=r\circ\lambda_{j}$.
    3. (c)
      $(\lambda_{i})_{i\in I}\in\Phi(\mathbb{F}_{\theta})$.
The first item can be deduced by applying the law $h\circ\mathsf{L}_{M}=\mathsf{L}_{M}\circ h$ for $h\in\operatorname{End}({}_{\mathcal{R}}V)$ and $M\in\mathcal{R}$, in particular for $M=(m_{ij})_{i,j\in I}$ where for some $k$ we have $m_{kk}=\mathrm{id}_{C_{k}}$ and $m_{ij}=O_{ij}$ for all $(i,j)\neq(k,k)$. A similar argument establishes the equivalence of (2a) and (2b).
It remains to prove (2b) $\Leftrightarrow$ (2c). First we record a helpful observation.
Claim 1. For all $i,j\in I$ and $r\in R_{ij}$, there exists a term $t(x,y_{1},\ldots,y_{n})$ and $\alpha$-classes $E_{1},\ldots,E_{n}$ such that for all $a\in C_{j}$,
$(r(a),e_{i})/\Delta=t^{D({\mathbf{A}},\theta)}((a,e_{j})/\Delta,0_{E_{1}},\ldots,0_{E_{n}}).$
To prove Claim 1, choose $f\in\operatorname{Pol}_{1}({\mathbf{A}})$ with $f(e_{j})=e_{i}$ and $r=f{\upharpoonright}_{C_{j}}$, and pick a term $t(x,y_{1},\ldots,y_{n})$ and $\mathbf{u}\in A^{n}$ so that $f(x)=t(x,\mathbf{u})$. Then for each $\ell\in[n]$ let $E_{\ell}=u_{\ell}/\alpha$. Then for any $a\in C_{j}$,
$(r(a),e_{i})/\Delta$ $=(t(a,\mathbf{u}),t(e_{j},\mathbf{u}))/\Delta$
$=t^{D({\mathbf{A}},\theta)}((a,e_{j})/\Delta,(u_{1},u_{1})/\Delta,\ldots,(u_{n},u_{n})/\Delta)$
$=t^{D({\mathbf{A}},\theta)}((a,e_{j})/\Delta,0_{E_{1}},\ldots,0_{E_{n}}),$
which proves Claim 1.
Now assume that $(\lambda_{i})_{i\in I}=\Phi(\lambda)$ with $\lambda\in\mathbb{F}_{\theta}$. Fix $i,j\in I$ and $r\in R_{ij}$ and let $a\in C_{j}$. We must show $\lambda_{i}(r(a))=r(\lambda_{j}(a))$. Choose $t(x,\mathbf{y})$ and $E_{1},\ldots,E_{n}\in A/\alpha$ for $i,j,r$ as in Claim 1. By definition, $\lambda_{i}(r(a))$ is the unique $b\in C_{i}$ such that $\lambda((r(a),e_{i})/\Delta)=(b,e_{i})/\Delta$. Well,
$\lambda((r(a),e_{i})/\Delta)$ $=\lambda(t^{D({\mathbf{A}},\theta)}((a,e_{j})/\Delta,0_{E_{1}},\ldots,0_{E_{n}}))$
$=t^{D({\mathbf{A}},\theta)}(\lambda((a,e_{j})/\Delta),0_{E_{1}},\ldots,0_{E_{n}})$ as $\lambda\in\mathbb{F}_{\theta}$
$=t^{D({\mathbf{A}},\theta)}((\lambda_{j}(a),e_{j})/\Delta,0_{E_{1}},\ldots,0_{E_{n}})$
$=(r(\lambda_{j}(a)),e_{i})/\Delta,$
which implies $\lambda_{i}(r(a))=r(\lambda_{j}(a))$, proving (2c) $\Rightarrow$ (2b).
For the opposite implication, assume that $(\lambda_{i})_{i\in I}$ satisfies the condition of (2b). We first claim that for all $i,j\in I$, all $a\in C_{i}$, and all $b\in C_{j}$,
(8.1) $(a,e_{i})\stackrel{{\scriptstyle\Delta}}{{\equiv}}(b,e_{j})~~\mbox{implies}~~(\lambda_{i}(a),e_{i})\stackrel{{\scriptstyle\Delta}}{{\equiv}}(\lambda_{j}(b),e_{j}).$
Indeed, assume $(a,e_{i})\stackrel{{\scriptstyle\Delta}}{{\equiv}}(b,e_{j})$. Choose a $\theta$-class $C_{k}$ so that $C_{i},C_{j}\rightarrow C_{k}$ (see Lemma 5.7). Let $f_{i},f_{j}\in\operatorname{Pol}_{1}({\mathbf{A}})$ be witnesses to $C_{i}\rightarrow C_{k}$ and $C_{j}\rightarrow C_{k}$ respectively. By Lemma 5.7, we can assume $f_{i},f_{j}$ are chosen so that $f_{\ell}(e_{\ell})=e_{k}$ for $\ell=i,j$. Then $(f_{i}(a),e_{k})\stackrel{{\scriptstyle\Delta}}{{\equiv}}(a,e_{i})\stackrel{{\scriptstyle\Delta}}{{\equiv}}(b,e_{j})\stackrel{{\scriptstyle\Delta}}{{\equiv}}(f_{j}(b),e_{k})$ by Lemma 5.7(1) and the assumption, which implies $f_{i}(a)=f_{j}(b)$ by Lemma 5.2(1). Let $r=f_{i}{\upharpoonright}_{\theta_{i}}$ and $s=f_{j}{\upharpoonright}_{\theta_{j}}$ and note that $r\in R_{ki}$ and $s\in R_{kj}$. Condition (2b) then gives $\lambda_{k}\circ r=r\circ\lambda_{i}$ and $\lambda_{k}\circ s=s\circ\lambda_{j}$. Then
$(\lambda_{i}(a),e_{i})\stackrel{{\scriptstyle\Delta}}{{\equiv}}(r(\lambda_{i}(a)),e_{k})$ $=(\lambda_{k}(r(a)),e_{k})$
$=(\lambda_{k}(s(b)),e_{k})=(s(\lambda_{j}(b)),e_{k})\stackrel{{\scriptstyle\Delta}}{{\equiv}}(\lambda_{j}(b),e_{j})$
which proves (8.1).
Now let $\lambda:D({\mathbf{A}},\theta)\rightarrow D({\mathbf{A}},\theta)$ be the unique map satisfying
$\lambda((x,e_{i})/\Delta)=(\lambda_{i}(x),e_{i})/\Delta$
for all $i\in I$ and all $x\in C_{i}$. ($\lambda$ is well-defined by (8.1).) We will show that $\lambda\in\mathbb{F}_{\theta}$, which will complete the proof of Proposition A1.1. Clearly $\lambda(\varphi)\subseteq\varphi$ and $\lambda(0_{E})=0_{E}$ for all $E\in A/\alpha$. It remains to show that $\lambda\in\operatorname{End}(D({\mathbf{A}},\theta))$. The next Claim is an immediate consequence of Proposition 3.8.
Claim 2. For all $n\geq 1$, all $f\in\operatorname{Pol}_{n}({\mathbf{A}})$, and all $\theta$-classes $C_{i_{1}},\ldots,C_{i_{n}}$, if $C_{\ell}$ is the $\theta$-class containing $f(e_{i_{1}},\ldots,e_{i_{n}})$, then there exist $r_{t}\in R_{\ell i_{t}}$ for $t\in[n]$ such that the following holds: for all $(a_{1},\ldots,a_{n})\in C_{i_{1}}\times\cdots\times C_{i_{n}}$,
[…]
Now let $F$ be an $n$-ary basic operation. Choose $n$ elements of $D({\mathbf{A}},\theta)$, which may be written $(a_{1},e_{i_{1}})/\Delta,\ldots,(a_{n},e_{i_{n}})/\Delta$ for some $\theta$-classes $\theta_{i_{t}}$ and $a_{t}\in C_{i_{t}}$. We must show
(8.2) $\lambda(F((a_{1},e_{i_{1}})/\Delta,\ldots,(a_{n},e_{i_{n}})/\Delta))=F(\lambda((a_{1},e_{i_{1}})/\Delta),\ldots,\lambda((a_{n},e_{i_{n}})/\Delta)).$
Let $C_{\ell}$ be the $\theta$-class containing $F(\mathbf{a})$ and choose $r_{t}\in R_{\ell i_{t}}$ for $t\in[n]$ as in Claim 2 for $F$ and $C_{i_{1}},\ldots,C_{i_{n}}$. We have $(F(\mathbf{a}),F(\mathbf{e}))\stackrel{{\scriptstyle\Delta}}{{\equiv}}(d(F(\mathbf{a}),F(\mathbf{e}),e_{\ell}),e_{\ell})$ by Lemma 5.3(1). Thus the left-hand side of (8.2) can be written
$\lambda(F((a_{1},e_{i_{1}})/\Delta,\ldots,(a_{n},e_{i_{n}})/\Delta))$ $=\lambda((F(\mathbf{a}),F(\mathbf{e}))/\Delta)$
$=\lambda((d(F(\mathbf{a}),F(\mathbf{e}),e_{\ell}))/\Delta)$
$=(\lambda_{\ell}(d(F(\mathbf{a}),F(\mathbf{e}),e_{\ell})),e_{\ell})/\Delta.$
The right-hand side of (8.2) can be written
$F(\lambda((a_{1},e_{i_{1}})/\Delta),\ldots,\lambda((a_{n},e_{i_{n}})/\Delta))$ $=F((\lambda_{i_{1}}(a_{1}),e_{i_{1}})/\Delta,\ldots,(\lambda_{i_{n}}(a_{n}),e_{i_{n}})/\Delta)$
$=(F(\lambda_{i_{1}}(a_{1}),\ldots,\lambda_{i_{n}}(a_{n})),F(\mathbf{e}))/\Delta$
$=(d(F(\lambda_{i-1}(a_{1}),\ldots,\lambda_{i_{n}}(a_{n})),F(\mathbf{e}),e_{\ell}),e_{\ell})/\Delta.$
Thus to prove (8.2), it will suffice to show
(8.3) $\lambda_{\ell}(d(F(\mathbf{a}),F(\mathbf{e}),e_{\ell}))=d(F(\lambda_{i_{1}}(a_{1}),\ldots,\lambda_{i_{n}}(a_{n})),F(\mathbf{e}),e_{\ell}).$
By Claim 2, there exist $r_{t}\in R_{\ell i_{t}}$ for $t\in[n]$ such that for all $(x_{1},\ldots,x_{n})\in C_{i_{1}}\times\cdots\times C_{i_{n}}$, $F(\mathbf{x})=\sum_{t=1}^{n}r_{t}(x_{t})+F(\mathbf{e})$. Recall that each $\lambda_{j}$ and $r_{t}$ is a group homomorphism between the appropriate groups from $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e_{i_{1}}),\ldots,\operatorname{Grp}_{{\mathbf{A}}}(\theta,e_{i_{n}}),\operatorname{Grp}_{{\mathbf{A}}}(\theta,e_{\ell})$. Thus computing in $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e_{\ell})$,
$\lambda_{\ell}(d(F(\mathbf{a}),F(\mathbf{e}),e_{\ell}))$ $=\lambda_{\ell}(F(\mathbf{d}))-\lambda_{\ell}(F(\mathbf{e}))$
$=\lambda_{\ell}\left(\sum_{t=1}^{n}r_{t}(a_{t})+F(\mathbf{e})\right)-\lambda_{\ell}(F(\mathbf{e}))$
$=\sum_{t=1}^{n}\lambda_{\ell}(r_{t}(a_{t}))$ $(\dagger)$
while
$d(F(\lambda_{i_{1}}(a_{1}),\ldots,\lambda_{i_{n}}(a_{n})),F(\mathbf{e}),e_{\ell})$ $=\left(\sum_{t=1}^{n}r_{t}(\lambda_{i_{t}}(a_{t}))+F(\mathbf{e})\right)-F(\mathbf{e})$
$=\sum_{t=1}^{n}r_{t}(\lambda_{i_{t}}(a_{t}).$ $(\ddagger)$
Since $\lambda_{\ell}\circ r_{t}=r_{t}\circ\lambda_{i_{t}}$ for each $t\in[n]$ by assumption (2b), we have established (8.3) and so have finished the proof of Proposition A1.1. ∎
Finally, we show that $\mathbb{F}_{\theta}$ coincides with the finite field $\mathbb{F}$ constructed in the proof of Theorem 3.9 when $\mathbf{A}$} is finite.
Proposition A1.2.
Suppose $\mathcal{V}$ is a variety with a weak difference term $d$, $\mathbf{A}$} is a finite algebra in $\mathcal{V}$, and $\theta$ is an abelian minimal congruence of $\mathbf{A}$}. Let $\mathbb{F}$ be the finite field associated to $\theta$ in the proof of Theorem 3.9. That is, let $C_{1},\ldots,C_{k}$ be a list of the $\theta$-classes, let $\overline{C}:=C_{1}\times\cdots\times C_{k}$, choose $\mathbf{e}=(e_{1},\ldots,e_{k})\in\overline{C}$, and define ++ on $\overline{C}$ by $x+y:=d(x,\mathbf{e},y)$ so that ${\mathbf{C}}:=(\overline{C},+,\mathbf{e})$ is an abelian group. Let $\mathfrak{C}({\mathbf{A}},\nu_{\theta})$ be defined as in the proof of Theorem 3.9, let $\mathbf{R}$} be the subring of $\operatorname{End}({\mathbf{C}})$ such that $\mathfrak{C}({\mathbf{A}},\nu_{\theta})$ is polynomially equivalent to $_{{\mathbf{R}}}\hskip-1.0pt{\mathbf{C}}$, and let $\mathbb{F}=\operatorname{End}({}_{{\mathbf{R}}}\hskip-1.0pt{\mathbf{C}})$. Then $\mathbb{F}\cong\mathbb{F}_{\theta}$.
Proof sketch.
First recall the construction of Freese’s division ring $\mathbb{D}$ associated to $\theta$ and the transversal $T:=\{e_{1},\ldots,e_{k}\}$ described at the beginning of this section. By finiteness, the abelian group $V$ in Freese’s construction coincides with the group $\mathbf{C}$}. Also recall the unital subring $\mathsf{L}_{\mathcal{R}}$ of $\mathbf{End}({\mathbf{C}})$ from the construction of $\mathbb{D}$. If we can prove $\mathsf{L}_{\mathcal{R}}={\mathbf{R}}$, then it will follow that $\mathbb{D}=\mathbb{F}$ and therefore $\mathbb{F}\cong\mathbb{F}_{\theta}$ by Proposition A1.1.
Following [8], the universe of $\mathbf{R}$} may be given by
$R=\{f(x)\in\operatorname{Pol}_{1}(\mathfrak{C}({\mathbf{A}},\nu_{\theta}))\,:\,\text{$f(\mathbf{e})=\mathbf{e}$}\}.$
By [23, Remark 2.8], the unary polynomials of $\mathfrak{C}({\mathbf{A}},\nu_{\theta})$ are the functions $f:C\rightarrow C$ which can be described as follows: for some $k$-ary polynomials $f_{1},\ldots,f_{k}\in\operatorname{Pol}_{k}({\mathbf{A}})$ with $f_{i}(\mathbf{e})\in C_{i}$ for $i\in[k]$,
[…]
Assuming $f\in R$, i.e., $f(\mathbf{e})=\mathbf{e}$, we get $f_{i}(\mathbf{e})=e_{i}$ for each $i\in[k]$. Thus for each $i\in[k]$ there exist $r_{1},\ldots,r_{k}\in R_{ij}$ (see Freese’s construction) so that for all $\mathbf{c}=(c_{1},\ldots,c_{k})\in C$,
$f_{i}(\mathbf{c})=\sum_{j=1}^{k}r_{ij}(c_{j})\quad\mbox{computed in $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e_{i})$},$
by Proposition 3.8. That is, $f=\mathsf{L}_{M}$ where $M=(r_{ij})_{i,j\in[k]}\in\mathcal{R}$ (assuming $f\in R$). The logic is reversible and we have $R=\mathsf{L}_{\mathcal{R}}$ as desired. ∎
Appendix 2: Examples
In this appendix we illustrate some of our results from Sections 5 and 6 by presenting a recipe for constructing certain finite subdirectly irreducible algebras with abelian monoliths in varieties having weak difference terms. The algebras $\mathbf{A}$} can be constructed so that if $\mu$ is the monolith, $\alpha=(0:\mu)$, and $\varphi=\overline{\alpha}/\Delta_{\mu,\alpha}$, then
  1. (1)
    $\mathbb{F}_{\mu}$ can be any desired finite field $\mathbb{F}$;
@@
    The vector spaces $_{\mathbb{F}}\hskip-2.0pt\operatorname{Grp}_{D({\mathbf{A}})}(\varphi,0_{E})$ as $E$ ranges over the $\alpha$-classes can have any desired (finite) dimensions.
@@
    The ranges of the $\mu$-classes within a given $\alpha$-class $E$ can be any collection of subspaces of $_{\mathbb{F}}\hskip-2.0pt\operatorname{Grp}_{D({\mathbf{A}})}(\varphi,0_{E})$, as long as the collection includes $0_{E}/\varphi$ itself.
Our recipe is a variant of a known construction [3]. We first give a useful method for constructing varieties with a prescribed weak difference term. For this, we need the following result characterizing weak difference terms at the level of varieties. Following [18], given a variety $\mathcal{V}$, let ${\mathbf{F}}={\mathbf{F}}_{\mathcal{V}}(3)$ be the $\mathcal{V}$-free algebra freely generated by $\{x,y,z\}$, let $\alpha=\operatorname{Cg}(x,z)$, $\beta=\operatorname{Cg}(x,y)$ and $\gamma=\operatorname{Cg}(y,z)$ calculated in $\mathbf{F}$}, and define $\beta_{0}=\beta$, $\gamma_{0}=\gamma$, $\beta_{n+1}=\beta\vee(\alpha\wedge\gamma_{n})$, and $\gamma_{n+1}=\gamma\vee(\alpha\wedge\beta_{n})$ for $n\geq 0$.
Theorem A2.1.
@@
(1) $\Rightarrow$ (3) can be extracted from the proof of Theorem 4.8 (1) $\Rightarrow$ (2) $\Rightarrow$ (3) in [18]. Standard arguments with roots in [27, 32] can produce sets $\Sigma_{0},\Sigma_{1},\Sigma_{2},\ldots$ of identities witnessing (2) $\Leftrightarrow$ (3). To prove (3) $\Rightarrow$ (1), assume that $(x,d)\in\gamma_{m}$ and $(d,z)\in\beta_{m}$. Then by [33, Theorem 2.1 (3) $\Rightarrow$ (4)], there exist idempotent ternary terms $f_{1},\ldots,f_{n},g_{1},\ldots,g_{n}$ (for some $n$) such that $\mathcal{V}\models f_{i}(x,y,x)\approx g_{i}(x,y,x)$ for each $i\in[n]$, and for all ${\mathbf{A}}\in\mathcal{V}$ and all $a,b\in A$, if
@@
then $d(a,a,b)=b=d(b,a,a)$ and $d(b,b,a)=a=d(a,b,b)$. Now one can easily adapt the proof of [13, Theorem 1.2 (3) $\Rightarrow$ (1)] to get that $d$ is a weak difference term for $\mathcal{V}$. ∎
Definition A2.2.
A semilattice-over-Maltsev operation is any operation $d:A^{3}\rightarrow A$ constructed in the following way. Let $(S,\wedge)$ be a (meet) semilattice. For each $s\in S$ let $V_{s}$ be a non-empty set, and assume that $V_{s}\cap V_{t}=\varnothing$ for all $s,t\in S$ with $s\neq t$. For each pair $s,t\in S$ with $s\geq t$, assign a function $f_{(s,t)}:V_{s}\rightarrow V_{t}$, so that $f_{(s,s)}=\mathrm{id}_{V_{s}}$. Finally, for each $s\in S$ let $m_{s}:(V_{s})^{3}\rightarrow V_{s}$ be a Maltsev operation.
@@
The identities (8.5) and (8.6) therefore hold in $\mathsf{HSP}({\mathbf{A}})$. Let ${\mathbf{F}}={\mathbf{F}}_{\mathsf{HSP}({\mathbf{A}})}(3)$ be the free algebra in $\mathsf{HSP}({\mathbf{A}})$ freely generated by $\{x,y,z\}$. As in the discussion preceding Theorem A2.1, let $\alpha=\operatorname{Cg}(x,z)$, $\beta=\operatorname{Cg}(x,y)$ and $\gamma=\operatorname{Cg}(y,z)$ be calculated in $\mathbf{F}$}. Interpreting $d$ and $f_{0},\ldots,f_{3},g_{0},\ldots,g_{3}$ as elements of $\mathbf{F}$} in the usual way, we find that the identities (8.5) and (8.6) insure that $f_{0},\ldots,f_{3},g_{0},\ldots,g_{3}$ witness
@@
Thus $d$ is a weak difference term for $\mathsf{HSP}({\mathbf{A}})$ by Theorem A2.1. ∎
@@
Choose vector spaces $V_{(\ell,i)}$ over $\mathbb{F}$ and isomorphisms $\sigma_{(\ell,i)}:V_{(\ell,i)}\cong W_{(\ell,i)}$ for $\ell\leq m$ and $1\leq i\leq n_{\ell}$, so that the spaces $V_{(\ell,i)}$ ($0\leq\ell\leq m$, $0\leq i\leq n_{\ell}$) are pairwise disjoint. Also set $\sigma_{(\ell,0)}=\mathrm{id}_{V_{(\ell,0)}}$ for each $\ell\leq m$. Let $E_{\ell}$ be the (disjoint) union of $V_{(\ell,0)},V_{(\ell,1)},\ldots,V_{(\ell,n_{\ell})}$, and let $A$ be the (disjoint) union of $E_{0},\ldots,E_{m}$. Then let $\mu$ and $\alpha$ be the equivalence relations on $A$ whose equivalence classes are the spaces $V_{(\ell,i)}$ ($\ell\leq m$, $i\leq n_{\ell}$) and the sets $E_{0},\ldots,E_{m}$ respectively. Also let $B=V_{(0,0)}\cup\cdots\cup V_{(m,0)}$ and let $\psi$ be the equivalence relation on $B$ whose equivalence classes are $V_{(0,0)},\ldots,V_{(m,0)}$. We will build an algebra $\mathbf{A}$} with universe $A$ such that:
@@
  • Letting $\Delta=\Delta_{\mu,\alpha}$ and $\varphi=\overline{\alpha}/\Delta$, there is a bijection from $D({\mathbf{A}},\mu)$ to $B$ which sends $\varphi$ to $\psi$ and sends each $\varphi$-class $(\mu\cap(E_{\ell})^{2})/\Delta$ to $V_{(\ell,0)}$.
  • Modulo this bijection, for each $\alpha$-class $E_{\ell}$, the ranges (see Definition 5.4) of the $\mu$-classes $V_{(\ell,i)}$ contained in $E_{\ell}$ correspond precisely to the subspaces $W_{(\ell,i)}$ in $V_{(\ell,0)}$.
  • The division ring $\mathbb{F}_{\mu}$ associated to $\mu$ is $\mathbb{F}$ (up to isomorphism).
@@
[…]
@@
Now let $d:A^{3}\rightarrow A$ be the semilattice-over-Maltsev operation (see Definition A2.2) defined by the data $(S,\wedge)$, $(V_{s}:s\in S)$, $(f_{(s,t)}:s,t\in S,\,s\geq t)$, and the family of Maltsev operations $m_{s}(x,y,z)=x-y+z$ evaluated in each vector space $V_{s}$. That is, for all $a\in V_{s_{1}}$, $b\in V_{s_{2}}$ and $c\in V_{s_{3}}$, letting $t=s_{1}\wedge s_{2}\wedge s_{3}$ we have
$\mbox{$d(a,b,c)=f_{(s_{1},t)}(a)-f_{(s_{2},t)}(b)+f_{(s_{3},t)}(c)$\quad calculated in $V_{t}$}.$
For each $\ell\leq m$ define $\sigma_{\ell}:E_{\ell}\rightarrow V_{(\ell,0)}$ so that $\sigma_{\ell}{\upharpoonright}_{V_{(\ell,i)}}=\sigma_{(\ell,i)}$ for all $i\leq n_{\ell}$. For each $s\in S$ let $0_{s}$ denote the zero element of $V_{s}$. We define some unary operations on $A$.
@@
    $F_{(\ell,i)}^{g}(a)=\left\{\begin{array}[]{cl}(g\circ\sigma_{\ell})(a)&\mbox{if $x\in E_{\ell}$}\\ 0_{(k,0)}&\mbox{if $a\in E_{k}$ with $k\neq\ell$.}\end{array}\right$.
@@
    $G_{\ell}(a)=\left\{\begin{array}[]{cl}(g_{\ell}\circ\sigma_{0})(a)&\mbox{if $a\in E_{0}$}\\ 0_{(k,0)}&\mbox{if $a\in E_{k}$ with $k\neq 0$}\end{array}\right$.
@@
    $G_{\ell}^{\prime}(a)=\left\{\begin{array}[]{cl}(h_{\ell}\circ\sigma_{\ell})(a)&\mbox{if $a\in E_{\ell}$}\\ 0_{(k,0)}&\mbox{if $a\in E_{k}$ with $k\neq\ell$}\end{array}\right$.